Can someone please help me with this:
Lighthouse Tours provide seven-day guided tours along the New England coast. The company pays its guides a total of $100,000 per year. The average cost of supplies, lodging and food per customer is $500. The company expects a total of 500 customers during the period January - June, and a total of 1,500 customers from July through December. Lighthouse wants to earn $200 per customer. For promotional reasons the company desires to charge the same price throughout the year. Based on this information, what is the correct price per customer?(Round to the nearest dollar)
a. $550
b. $600
c. $750
d. $800

Answer by gonzo(654)

answer is 750.
here's how.
guides = 100,000
number of customers is 2000 (500 first half of year, 1500 second half of year)
cost per customer is 500
500 * 2000 = 1,000,000 total customer cost.
total cost is 1,100,000
profit per customer is 200
total profit is 200 * 2000 = 400,000
total revenue required is cost plus profit is 1,100,000 + 400,000 = 1,500,000
total revenue required of 1,500,000 divided by 2000 customer is 750 per customer.